# Highest-weight vectors and three-point functions in GKO coset decomposition

### v1

@@ 3.4 Norms of the highest weight vectors
Note that although the Theorem 3.7 gives formula for $u_{l}$ in terms of Wakimoto free field realization in for the norm formula above the Shapovalov form in Verma module is used.
@@
Notation 3.16.
The highest weight vectors $|\tilde{u}_{l}(\lambda)\rangle$ are defined by
$|\tilde{u}_{l}(\lambda)\rangle=\begin{cases}|u_{l}(\lambda)\rangle,l>0;\\ \frac{|u_{l}(\lambda)\rangle}{\left\|u_{l}(\lambda)\right\|^{2}},l<0.\end{cases}$ (3.45)
@@
For $2l\in\mathbb{Z}$ the norm of the vector of $u_{l}(\lambda)$ is
@@
Here we used the following notation for the product over integer triangle
@@
We will also use notation for the product over integer segment
@@
Below, we will give the computational proof of Theorem 3.14. However, there is also a representation-theoretic proof based on Kac-Kazhdan theorem (see Theorem 2.4). Let us illustrate this with the first non-trivial vector $u_{-1/2}(\lambda)$. The existence of the pole $\lambda$ comes from the fact that for $\lambda=0$ the module $\mathcal{M}_{\lambda,k}$ has a singular vector $f_{0}v_{\lambda,k}$. On the other hand, the zero $\lambda+1$ corresponds to the fact that the module $\mathcal{M}_{\lambda+1,k+1}$ on the right side of decomposition (3.1b) has a singular vector.
@@ 3.5 Calculation of the norms
Remark that the operator $J(z)$ can be described as the operator $I(z)$ dressed by the screening $S(t)$ (see formula 2.52)
@@
The proof is similar to Proposition 3.11. ∎
We will also need screening current conjugated by $e^{-g_{0}}$. Namely, we have
$S(t)=e^{-\sqrt{\frac{2}{\kappa}}\varphi(t)}\beta(t)=e^{-g_{0}}\tilde{S}(t)e^{g_{0}},$ (3.53)
where $\tilde{S}(t)$ is defined by formula
$\tilde{S}(t)=e^{-\sqrt{\frac{2}{\kappa}}\varphi(t)}(\beta(t)+e^{(1)}(t)).$ (3.54)
@@
We will denote by $X^{\dagger}$ the operator conjugated to an operator $X$ with respect to the Shapovalov form.
@@
We will compute the proportionality coefficient $C(\kappa,\lambda)$ in the proof of Theorem 3.23.
@@
For $l\leq 0$ we have the formula The action of operator $J(z)$ on the highest weight vectors has the form
@@
Equivalently we have the formula for the action of operator $e^{g_{0}}J(z)e^{-g_{0}}$ on the extremal vectors
$e^{g_{0}}J(z)e^{-g_{0}}v_{l}(\lambda{+}1)=z^{l-\frac{\lambda+3}{\kappa}}\left((-1)^{2(l+1/2)(l+1)}C(\kappa,\lambda)^{-1}\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}v_{l+\frac{1}{2}}(\lambda)+O(z)\right).$ (3.61)
@@
$\left\langle v_{l+\frac{1}{2}}(\lambda),\,e^{g_{0}}J(1)e^{-g_{0}}v_{l}(\lambda+1)\right\rangle(-1)^{2(l+1/2)(l+1)}C(\kappa,\lambda)^{-1}\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}.$ (3.63)
@@
$\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}=C(\kappa,\lambda)~{}B(-\frac{1}{\kappa},2l-\frac{\lambda+1}{\kappa}+1).$ (3.66)
Using values for $\|u_{0}(\lambda)\|^{2}$ and $\|u_{-1/2}(\lambda)\|^{2}$ given in Example 3.15 we find that $C(\kappa,\lambda)={B(-\frac{1}{\kappa};-\frac{\lambda+1}{\kappa}+1)}^{-1}$. ∎
@@
$u_{l}(\lambda)=e^{g_{0}^{\dagger}}v_{l}(\lambda).$ (3.67)
@@ 4.1 Main Theorem
$\widetilde{\mathcal{U}}_{\nu,n}(x,z)=\mathtt{Y}_{1}\otimes\mathtt{Y}_{k}(|\tilde{u}_{n}(\nu)\rangle;x,z),\qquad\mathcal{U}_{\nu,n}(x,z)=\mathtt{Y}_{1}\otimes\mathtt{Y}_{k}(|u_{n}(\nu)\rangle;x,z).$ (4.2)
@@
The following proposition is operator reformulation of the fact that $\tilde{u}_{n}(\nu)$ is a highest weight vector with respect to $\widehat{\mathfrak{sl}}(2)_{k+1}$ and coset Virasoro.
@@
$\tilde{C}_{m,n,l}(\mu,\nu,\lambda)=\langle\tilde{u}_{m}(\mu)|\widetilde{\mathcal{U}}_{\nu,n}(1,1)|\tilde{u}_{l}(\lambda)\rangle.$ (4.5)
It follows from (2.48) that $\tilde{C}_{m,n,l}$ vanishes unless $l+m+n\in\mathbb{Z}$. To be more precise, since normalization of $\mathcal{U}_{\nu}$ is not given, each individual $\tilde{C}_{m,n,l}$ is not defined. However, the ratio of $\widetilde{\mathcal{U}}_{\nu,n}$ that corresponds to the same representations pf $\widehat{\mathfrak{sl}}(2)_{1}\oplus\widehat{\mathfrak{sl}}(2)_{k}$ is well defined. For example one can ask for the ratio with the three-point functions of highest weight vectors
$\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\tilde{C}_{\{m\},\{n\},\{l\}}(\mu,\nu,\lambda)}$ (4.6)
@@
$\langle v_{\mu,k}|\mathcal{U}_{\nu}(1,1)|v_{\lambda,k}\rangle=\langle\tilde{u}_{0}(\mu)|\widetilde{\mathcal{U}}_{\nu,0}(1,1)|\tilde{u}_{0}(\lambda)\rangle=\langle\tilde{u}_{{1}/{2}}(\mu)|\widetilde{\mathcal{U}}_{\nu,0}(1,1)|\tilde{u}_{{1}/{2}}(\lambda)\rangle\\ =\langle\tilde{u}_{0}(\mu)|\widetilde{\mathcal{U}}_{\nu,1/2}(1,1)|\tilde{u}_{{1}/{2}}(\lambda)\rangle.=\langle\tilde{u}_{{1}/{2}}(\mu)|\widetilde{\mathcal{U}}_{\nu,1/2}(1,1)|\tilde{u}_{0}(\lambda)\rangle.$ (4.7)
@@
Observation 1 The three-point function satisfies recurrence relation on $l$. This relation is formulated in the Proposition 4.25. This relation is proven by considering a four-point conformal block with the insertion of a degenerate field.
@@
Consider particular case $n=-\frac{1}{2},\nu=1$. The corresponding operator is given by the formula 4.3c. Assume now that $\mu=\lambda+1$, then $\widetilde{\mathcal{U}}_{-1/2,1}$ coincides with $\frac{1}{2}I(z)$ and using formula (3.41) and Notation 3.16 we have
@@
This case corresponds to matrix elements of $b(x,z)$, see we calculate them during the proof of main theorem below, wee Proposition 4.13
@@
In the case $\lambda+\nu-\mu=2N$ or $\lambda-\nu-2-\mu=2N$ one can use bosonizations of vertex operator given on Propositions 2.27, 2.28. This leads to new Selberg-type integrals, see Theorem 6.3.
The remaining part of this section is organized as follows. In section 4.2 we study symmetry of three-point function and prove relations (4.9). Then we focus on the proof of Proposition 4.25.
The idea is to consider four point conformal block $\langle\tilde{u}_{m}(\mu)|\widetilde{\mathcal{U}}_{\nu,n}(1,1)b(x,z)|\tilde{u}_{l}(\lambda)\rangle$ with insertion of degenerate field $b(x,z)$. We study properties of operator $b(x,z)$ in Section 4.3. In Sections 4.4 and 4.5 we recall definitions of conformal blocks, BPZ and KZ equations, and their solutions in terms of hypergeometric functions. We put all things together in Section 4.6. Firstly we prove Proposition 4.25 using triviality of monodromy of conformal block with insertion of $b(x,z)$, this is a variation of an argument used in [ZZ89], [Tes95], [Tes99], [BS15]. Then we consider generic conformal blocks and use Theorem 4.5 to find conformal blocks relations which are equivalent to blowup relations.
4.2 Symmetries of three point functions
@@
$\mathtt{m}=\mathtt{m}_{1}\otimes\mathtt{m}_{k}:\big{(}\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}(0,0)\big{)}\otimes\big{(}\mathcal{L}_{j,1}\otimes\mathcal{M}_{\nu,k}(1,1)\big{)}\otimes\big{(}\mathcal{L}_{r,1}\otimes\mathcal{M}_{\mu,k}(\infty,\infty)\big{)}\rightarrow\mathbb{C}$ (4.14)
@@
By straightforward (however not very illuminating) case-by-case check on can see that this formula is equivalent to (4.9a).
@@
$\alpha^{(0)}_{x,z}\big{(}\tilde{u}_{n}(\nu)\big{)}=x^{-2\lfloor n\rfloor}(-z^{2})^{n^{2}-\{n\}^{2}}\alpha^{(\infty)}_{y,w}\big{(}\tilde{u}_{n}(\nu)\big{)}$ (4.18)
@@
$(-1)^{n^{2}-\{n\}^{2}+\lfloor l\rfloor+\lfloor n\rfloor-\lfloor m\rfloor}\frac{\tilde{C}_{l,n,m}(\lambda,\nu,\mu)}{\tilde{C}_{0,0,0}(\lambda,\nu,\mu)}=\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\tilde{C}_{0,0,0}(\mu,\nu,\lambda)}$ (4.19)
@@ 4.3 Matrix elements of $b(x,z)$
Consider the operator $b(x,z)$ acting $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}\rightarrow\overline{\mathcal{L}_{1-i,1}\otimes\mathcal{M}_{\mu,k}}$.
@@
The proof is similar to the proof of Proposition 3.11. This operator is obtained from is obtained via operator-state correspondence map $\mathtt{Y}\otimes\mathtt{Y}$ applied to the vector
@@
Clearly this vector is highest weigh vector for $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$. Then it remains to show that
@@
Nontrivial matrix elements of $b(x,z)$ are equal to norms of the highest weight vectors, more explicitly
@@
$(\lambda+2m)\left\langle\tilde{u}_{m}(\lambda),b_{0}(z)\,\tilde{u}_{l}(\lambda)\right\rangle=\left\langle\tilde{u}_{m}(\lambda),h_{0}^{\Delta}b_{0}(z)\,\tilde{u}_{l}(\lambda)\right\rangle=(\lambda+2l+1)\left\langle\tilde{u}_{m}(\lambda),b_{0}(z)\,\tilde{u}_{b}(\lambda)\right\rangle.$ (4.24)
@@
$\left\langle u_{-l+1/2}(\lambda),b_{0}(z)\,u_{-l}(\lambda)\right\rangle=\left\langle v_{-l+1/2}(\lambda),e^{-g_{0}^{\dagger}}e^{-g_{0}}b_{0}(z)\,v_{-l}(\lambda)\right\rangle=\\ =(-1)^{l(2l+1)}\left\langle v_{-l+1/2}(\lambda),e^{-g_{0}^{\dagger}}e^{-g_{0}}\,z^{-l}(v_{-l+1/2}(\lambda)+o(1))\right\rangle=(-1)^{l(2l+1)}z^{-l}\|u_{-l+1/2}\|^{2}.$ (4.26)
@@
$B^{-}_{m}(\mu)=\left\langle u_{m-1/2}(\mu),b_{1}(z)\,u_{m}(\mu)\right\rangle=\left\langle v_{m-1/2}(\mu),e^{g_{0}}b_{1}(z)e^{g_{0}^{\dagger}}\,v_{m}(\mu)\right\rangle\\ =\left\langle v_{m-1/2}(\mu),e^{g_{0}}e^{g_{0}^{\dagger}}b_{1}(z)\,v_{m+1/2}(\mu)\right\rangle=(-1)^{m(2m-1)}\|u_{m-1/2}(\mu)\|^{2}.$ (4.28)
@@
$(-1)^{l(2l-1)}\|u_{-l+1/2}(\lambda)\|^{2}=B^{+}_{-l}(\lambda)=\left\langle u_{-l+1/2}(\lambda)\big{|}b_{0}(1)\big{|}u_{-l}(\lambda)\right\rangle=\\ =\langle\tau\otimes\tau(u_{-l}(\lambda))|\tau\otimes\tau(b_{0}(1))|\tau\otimes\tau(u_{-l}(\lambda))\rangle=B^{-}_{l+1/2}(k-\lambda)=(-1)^{l(2l-1)}\|u_{l}(k-\lambda)\|^{2}.$ (4.30)

### v3

@@ 3.4 Norms of the highest weight vectors
Note that although the Theorem 3.7 provides a formula for $u_{l}$ in terms of the Wakimoto free field realization, the norm formula above uses the Shapovalov form in the Verma module.
@@
Notation 3.17.
The highest weight vectors $\tilde{u}_{l}(\lambda)$ are defined by
$\tilde{u}_{l}(\lambda)=\begin{cases}u_{l}(\lambda),l>0;\\ \frac{u_{l}(\lambda)}{\left\|u_{l}(\lambda)\right\|^{2}},l<0.\end{cases}$ (3.45)
@@
For $2l\in\mathbb{Z}$ the norm of the vector of $\tilde{u}_{l}(\lambda)$ is
@@
Here we used the following notation for the product over an integral triangle
@@
We will also use notation for the product over an integral segment
@@
Below, we will give the computational proof of Theorem 3.15. However, there is also a representation-theoretic proof based on Kac-Kazhdan theorem (see Theorem 2.4). Let us illustrate this with the first non-trivial vector $u_{-1/2}(\lambda)$. Its norm is equal to $(\lambda+1)/\lambda$, see formula (3.43), i.e., it has one zero and one pole.
The existence of the pole at $\lambda=0$ in $\left\|u_{-1/2}(\lambda)\right\|^{2}$ comes from the fact that for $\lambda=0$ the module $\mathcal{M}_{\lambda,k}$ has a singular vector $f_{0}v_{\lambda,k}$. Hence, the vector $v_{1/2}\otimes f_{0}v_{\lambda,k}\in\mathcal{L}_{1,1}\otimes\mathcal{M}_{\lambda,k}$ is the highest weight vector with zero norm. This vector does not agree with normalization (3.7), actually as one can see from the formula (3.8b) that the vector $u_{-1/2}(\lambda)$ has a pole at $\lambda=0$ with residue given by $v_{1/2}\otimes f_{0}v_{\lambda,k}$.
On the other hand, the zero $\lambda+1$ in $\left\|u_{-1/2}(\lambda)\right\|^{2}$ corresponds to the fact that the module $\mathcal{M}_{\lambda+1,k+1}$ on the right side of decomposition (3.1b) has a singular vector. This singular vector (with zero norm) is equal to $u_{-1/2}(\lambda)$. Namely, one can see from the formula (3.8b) that the vector $u_{-1/2}(\lambda)$ for $\lambda=-1$ is equal to $f_{0}^{\Delta}u_{1/2}(\lambda)$. This also implies that decomposition (3.1b) fails for this $\lambda$.
As a more subtle example let us consider the vector $u_{-1}(\lambda)$. Its expression (3.8) has a pole at $\lambda=0$, while the norm (3.43) has neither pole, nor zero at this point. Representation-theoretic reason for this is that, for $\lambda=0$, the modules on both left and right sides of decomposition (3.1a) has singular vectors. Namely, let $v_{-2}=f_{0}v_{0}\in\mathcal{L}_{0,k}$ be the singular vector on the left side. It generates a submodule isomorphic to $\mathcal{M}_{-2,k}$ and one can then study the submodule in the tensor product $\mathcal{L}_{0,1}\otimes\mathcal{M}_{-2,k}\subset\mathcal{L}_{0,1}\otimes\mathcal{M}_{0,k}$ bearing in mind the decomposition (3.1a). Then, the vector $u_{1}(-2)$ is a highest weight vector with highest weight $(0,k+1)$ and $f_{0}^{\Delta}u_{1}(-2)$ is a singular vector. The norm of $f_{0}^{\Delta}u_{1}(-2)$ should have a double zero at $\lambda=0$ (since both $f_{0}v_{0}$ and $f_{0}u_{1}(-2)$ have simple zeroes). On the other hand, this vector is proportional to the residue $\operatorname{Res}_{\lambda=0}u_{1}(\lambda)\sim f_{0}u_{1}(-2)$.
@@ 3.5 Calculation of the norms
Remark that the operator $J(z)$ can be described as the operator $I(z)$ dressed by the screening $S(t)$ (see formula (2.52))
@@
The same as proof of Proposition 3.12. ∎
We will also need screening current $S(t)=:\!{\mathrm{e}}^{-\sqrt{\frac{2}{\kappa}}\varphi(t)}\!\!:\beta(t)$ conjugated by ${\mathrm{e}}^{-g_{0}}$. Namely, we have
$\tilde{S}(t)={\mathrm{e}}^{g_{0}}S(t){\mathrm{e}}^{-g_{0}}=:\!{\mathrm{e}}^{-\sqrt{\frac{2}{\kappa}}\varphi(t)}\!\!:\big{(}\beta(t)-e^{(1)}(t)\big{)}.$ (3.53)
@@
Recall that $X^{\dagger}$ denotes the operator conjugated to an operator $X$ with respect to the Shapovalov form (see Definitions 3.3, 3.4).
@@
We will actually compute the proportionality coefficient $C(\kappa,\lambda)$ in the proof of Theorem 3.24.
@@
For $l\leq 0$ we have the formula for the action of operator $J(z)$ on the highest weight vectors
@@
Equivalently we have the formula for the action of operator ${\mathrm{e}}^{g_{0}}J(z){\mathrm{e}}^{-g_{0}}$ on the extremal vectors
${\mathrm{e}}^{g_{0}}J(z){\mathrm{e}}^{-g_{0}}v_{l}(\lambda{+}1)=z^{l-\frac{\lambda+3}{\kappa}}\left((-1)^{2(l+1/2)(l+1)}C(\kappa,\lambda)^{-1}\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}v_{l+\frac{1}{2}}(\lambda)+O(z)\right).$ (3.60)
@@
$\left\langle v_{l+\frac{1}{2}}(\lambda),\,{\mathrm{e}}^{g_{0}}J(1){\mathrm{e}}^{-g_{0}}v_{l}(\lambda+1)\right\rangle=(-1)^{2(l+1/2)(l+1)}C(\kappa,\lambda)^{-1}\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}.$ (3.62)
@@
$\frac{\|u_{l}(\lambda+1)\|^{2}}{\|u_{l+\frac{1}{2}}(\lambda)\|^{2}}=-C(\kappa,\lambda)B(-\frac{1}{\kappa},2l-\frac{\lambda+1}{\kappa}+1).$ (3.65)
Using values for $\|u_{0}(\lambda)\|^{2}$ and $\|u_{-1/2}(\lambda)\|^{2}$ given in Example 3.16 we find that $C(\kappa,\lambda)=-B(-\frac{1}{\kappa};-\frac{\lambda+1}{\kappa}+1)^{-1}$. ∎
@@
$u_{l}(\lambda)={\mathrm{e}}^{g_{0}^{\dagger}}v_{l}(\lambda).$ (3.66)
@@ 4.1 Main Theorem
$\widetilde{\mathcal{U}}_{\nu,n}(x,z)=\mathtt{Y}_{1}\otimes\mathtt{Y}_{k}\big{(}\tilde{u}_{n}(\nu);x,z\big{)},\qquad\mathcal{U}_{\nu,n}(x,z)=\mathtt{Y}_{1}\otimes\mathtt{Y}_{k}\big{(}u_{n}(\nu);x,z\big{)}.$ (4.2)
@@
The following proposition is an operator reformulation of the fact that $\tilde{u}_{n}(\nu)$ is a highest weight vector with respect to $\widehat{\mathfrak{sl}}(2)_{k+1}$ and coset Virasoro.
@@
$\tilde{C}_{m,n,l}(\mu,\nu,\lambda)=\big{\langle}\tilde{u}_{m}(\mu),\widetilde{\mathcal{U}}_{\nu,n}(1,1)\tilde{u}_{l}(\lambda)\big{\rangle}.$ (4.5)
It follows from fusion rules (2.48) that $\tilde{C}_{m,n,l}$ vanishes unless $l+m+n\in\mathbb{Z}$.
Note that, since normalization of $\mathcal{U}_{\nu}$ is not given, each individual $\tilde{C}_{m,n,l}$ is not defined. However, the ratios are well defined. For example, one can ask for the ratio with the three-point functions of highest weight vectors
$\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\tilde{C}_{\{m\},\{n\},\{l\}}(\mu,\nu,\lambda)}.$ (4.6)
@@
$\langle v_{\mu,k},\mathcal{U}_{\nu}(1,1)\,v_{\lambda,k}\rangle=\langle\tilde{u}_{0}(\mu)\ \widetilde{\mathcal{U}}_{\nu,0}(1,1)\,\tilde{u}_{0}(\lambda)\rangle=\langle\tilde{u}_{{1}/{2}}(\mu)\ \widetilde{\mathcal{U}}_{\nu,0}(1,1)\,\tilde{u}_{{1}/{2}}(\lambda)\rangle\\ =\langle\tilde{u}_{0}(\mu)\ \widetilde{\mathcal{U}}_{\nu,1/2}(1,1)\,\tilde{u}_{{1}/{2}}(\lambda)\rangle=\langle\tilde{u}_{{1}/{2}}(\mu)\ \widetilde{\mathcal{U}}_{\nu,1/2}(1,1)\,\tilde{u}_{0}(\lambda)\rangle.$ (4.7)
@@
Let us prove the theorem modulo the results to be proven later in this section.
@@
Observation 1 The three-point function satisfies recurrence relation on $l$. This relation is formulated in the Proposition 4.25. It is proven by considering a four-point conformal block with the insertion of a degenerate field.
@@
Consider particular case $n=-\frac{1}{2},\nu=1$. The corresponding operator is given by the formula (4.3c). Assume now that $\mu=\lambda+1$, then $\widetilde{\mathcal{U}}_{-1/2,1}$ coincides with $\frac{1}{2}I(z)$ and using formula (3.41) and Notation 3.17 we have
@@
This case corresponds to matrix elements of $b(x,z)$. We calculate them below during the proof of the main theorem, see Proposition 4.13.
@@
In the case $\lambda+\nu-\mu=2N$ or $\lambda-\nu-2-\mu=2N$ one can use bosonizations of the vertex operator given on Propositions 2.27, 2.28. This leads to new Selberg-type integrals, see Theorem 6.3.
The remaining part of this section is organized as follows. In section 4.2 we study symmetries of the three-point function and prove the relations (4.9). Then we focus on the proof of Proposition 4.25.
The idea is to consider four-point conformal block $\langle\tilde{u}_{m}(\mu),\widetilde{\mathcal{U}}_{\nu,n}(1,1)b(x,z)\tilde{u}_{l}(\lambda)\rangle$ with insertion of degenerate field $b(x,z)$. We study properties of operator $b(x,z)$ in Section 4.3. In Sections 4.4 and 4.5 we recall definitions of conformal blocks, BPZ and KZ equations, and their solutions in terms of hypergeometric functions. We put all things together in Section 4.6. Firstly we prove Proposition 4.25 using triviality of monodromy of the conformal block with insertion of $b(x,z)$, this is a variation of an argument used in [ZZ89], [Tes95], [Tes99], [BS15]. Then we consider generic conformal blocks and use Theorem 4.5 to find conformal blocks relations which are equivalent to the blowup relations.
4.2 Symmetries of three-point functions
@@
$\mathtt{m}=\mathtt{m}_{1}\otimes\mathtt{m}_{k}:\big{(}\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}(0,0)\big{)}\otimes\big{(}\mathcal{L}_{j,1}\otimes\mathcal{M}_{\nu,k}(1,1)\big{)}\otimes\big{(}\mathcal{L}_{r,1}\otimes\mathcal{M}_{\mu,k}(\infty,\infty)\big{)}\rightarrow\mathbb{C}.$ (4.14)
@@
By straightforward (and not illuminating) case-by-case check on can see that this formula is equivalent to (4.9a).
@@
$\alpha^{(0)}_{x,z}\big{(}\tilde{u}_{n}(\nu)\big{)}=x^{-2\lfloor n\rfloor}(-z^{2})^{n^{2}-\{n\}^{2}}\alpha^{(\infty)}_{y,w}\big{(}\tilde{u}_{n}(\nu)\big{)}.$ (4.18)
@@
$(-1)^{n^{2}-\{n\}^{2}+\lfloor l\rfloor+\lfloor n\rfloor-\lfloor m\rfloor}\frac{\tilde{C}_{l,n,m}(\lambda,\nu,\mu)}{\tilde{C}_{0,0,0}(\lambda,\nu,\mu)}=\frac{\tilde{C}_{m,n,l}(\mu,\nu,\lambda)}{\tilde{C}_{0,0,0}(\mu,\nu,\lambda)}.$ (4.19)
@@ 4.3 Matrix elements of $b(x,z)$
Recall the level 1 vertex operator $b(x,z)$ defined in sec. 2.6.3. We can consider it as an operator acting $\mathcal{L}_{i,1}\otimes\mathcal{M}_{\lambda,k}\rightarrow\overline{\mathcal{L}_{1-i,1}\otimes\mathcal{M}_{\lambda,k}}$. Taking into account decompositions (3.1) we can ask for the $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$ description.
@@
The proof is similar to the proof of Proposition 3.12. This operator is obtained by operator-state correspondence map $\mathtt{Y}\otimes\mathtt{Y}$ applied to the vector
@@
Clearly this vector is the highest weight vector for $\mathrm{Vir}^{\mathrm{coset}}\oplus\widehat{\mathfrak{sl}}(2)_{k+1}$. Then it remains to show that
@@
Nontrivial matrix elements of $b(x,z)$ are equal to norms of the highest weight vectors, namely
@@
$(\lambda+2m)\left\langle\tilde{u}_{m}(\lambda),b_{0}(z)\,\tilde{u}_{l}(\lambda)\right\rangle=\left\langle\tilde{u}_{m}(\lambda),h_{0}^{\Delta}b_{0}(z)\,\tilde{u}_{l}(\lambda)\right\rangle=(\lambda+2l+1)\left\langle\tilde{u}_{m}(\lambda),b_{0}(z)\,\tilde{u}_{l}(\lambda)\right\rangle.$ (4.24)
@@
$\left\langle u_{-l+1/2}(\lambda),b_{0}(z)\,u_{-l}(\lambda)\right\rangle=\left\langle v_{-l+1/2}(\lambda),{\mathrm{e}}^{-g_{0}^{\dagger}}{\mathrm{e}}^{-g_{0}}b_{0}(z)\,v_{-l}(\lambda)\right\rangle=\\ =(-1)^{l(2l+1)}\left\langle v_{-l+1/2}(\lambda),{\mathrm{e}}^{-g_{0}^{\dagger}}{\mathrm{e}}^{-g_{0}}\,z^{-l}(v_{-l+1/2}(\lambda)+o(1))\right\rangle=(-1)^{l(2l+1)}z^{-l}\|u_{-l+1/2}\|^{2}.$ (4.26)
@@
$B^{-}_{m}(\mu)=\left\langle u_{m-1/2}(\mu),b_{1}(z)\,u_{m}(\mu)\right\rangle=\left\langle v_{m-1/2}(\mu),{\mathrm{e}}^{g_{0}}b_{1}(z){\mathrm{e}}^{g_{0}^{\dagger}}\,v_{m}(\mu)\right\rangle\\ =\left\langle v_{m-1/2}(\mu),{\mathrm{e}}^{g_{0}}{\mathrm{e}}^{g_{0}^{\dagger}}b_{1}(z)\,v_{m}(\mu)\right\rangle=(-1)^{m(2m-1)}\|u_{m-1/2}(\mu)\|^{2}.$ (4.28)
@@
$(-1)^{l(2l-1)}\|u_{-l+1/2}(\lambda)\|^{2}=B^{+}_{-l}(\lambda)=\big{\langle}u_{-l+1/2}(\lambda),b_{0}(1)\,u_{-l}(\lambda)\big{\rangle}=\\ =\big{\langle}\tau\otimes\tau(u_{-l}(\lambda)),\tau\otimes\tau(b_{0}(1))\,\tau\otimes\tau(u_{-l}(\lambda))\big{\rangle}=B^{-}_{l+1/2}(k-\lambda)=(-1)^{l(2l-1)}\|u_{l}(k-\lambda)\|^{2}.$ (4.30)